which formula represents the interest I earned after the first month of an account that compounds monthly? p= principal r= interest rate
a. I = p*R
b. p=I*r /12
c I = P*R /12
d P=R*I /12

Answer by stanbon(57984)

which formula represents the interest I earned after the first month of an account that compounds monthly? p= principal r= interest rate
-----
Ans: I = P(R/12)
=======================================
Cheers,
Stan H.
===================

a. I = p*R
b. p=I*r /12
c I = P*R /12
d P=R*I /12